OpenStudy (anonymous):

Solve the equation -2sin^2 [theta] + cos [theta] = -2. Give exact solutions.

OpenStudy (anonymous):

I just would like some help with figuring this out. If someone can walk me through it, great. I don't want just an answer. I need to go through the steps.

OpenStudy (imstuck):

I'm looking for an identity now to help us solve this. Give me one sec, and I'll be back!

OpenStudy (imstuck):

Ok, here's what I came up with. If this was a "regular" polynomial, you would solve it by factoring, right? Well the same thing applies here, but you need to have either all sines or all cosines to do that. That's where our identities come in. I just wanted to make sure I had the right one before proceeding here. Are you familiar with the Pythagorean Identities in Trig?

OpenStudy (imstuck):

Are you still there?

OpenStudy (anonymous):

Oh geez, this could help me. Yes, I'm here.

OpenStudy (fibonaccichick666):

Do you know any identities that include \(sin^2(\theta)\) @WeAreSpoons ?

OpenStudy (anonymous):

Uh, so like what kind of things would that include?

OpenStudy (fibonaccichick666):

Go to the last chapter you learned about, what are the trig identities listed there?

OpenStudy (anonymous):

I'm sorry I sound like an idiot. I do want to learn and I learn fast. If you can take me through this, I will be very grateful.

OpenStudy (fibonaccichick666):

ok, take a look here: http://www.purplemath.com/modules/idents.htm

OpenStudy (anonymous):

okay so does this have anything to do with the half-angle formula? That includes sin^2 [theta] and it says that it equals 1 - cos (2[theta]) / 2

OpenStudy (fibonaccichick666):

Do you have a half angle?

OpenStudy (anonymous):

like here let me get a picture of what I mean

OpenStudy (anonymous):

where I suppose that (u) in the picture is the same as theta

OpenStudy (fibonaccichick666):

but we don't have a half angle, that won't help us

OpenStudy (anonymous):

ohhhhhhhhhhhh okay. is there another identity that includes sin^2 theta? I'll go look

OpenStudy (fibonaccichick666):

yup there is, it's usually the first one you learn

OpenStudy (anonymous):

is it a pythagorean identity?

OpenStudy (fibonaccichick666):

Just read, you shall see it

OpenStudy (fibonaccichick666):

It practically has strobe lights dancing off of it

OpenStudy (anonymous):

I feel so stupid.

OpenStudy (anonymous):

I don't even know where to start to fill in the holes in what I don't know. I'm in Pre-Calc. There's no book btw

OpenStudy (fibonaccichick666):

Don't. Tell me, which eq on that page have a sin^2 term

OpenStudy (fibonaccichick666):

NO BOOK??? What is wrong with that teacher...

OpenStudy (fibonaccichick666):

I shall get you a book

OpenStudy (fibonaccichick666):

http://home.comcast.net/~cherylwilcox1/site/

OpenStudy (anonymous):

the basic/pythagorean and the half-angle idents are the only ones I see with a sin^2 term

OpenStudy (fibonaccichick666):

Ok tell me each of them

OpenStudy (fibonaccichick666):

type them out here

OpenStudy (anonymous):

sin^2(t) + cos^2(t) = 1 That's one

OpenStudy (fibonaccichick666):

ok

OpenStudy (anonymous):

is that the one?

OpenStudy (fibonaccichick666):

What's the other option first

OpenStudy (anonymous):

uh, the half angle one, but you mentioned that there is no half angle here

OpenStudy (fibonaccichick666):

So what are your options?

OpenStudy (fibonaccichick666):

\[-2sin^2 [\theta] + cos [\theta] = -2\] Your eq.

OpenStudy (fibonaccichick666):

Do you have any options?

OpenStudy (anonymous):

This is the whole problem. I don't know if it includes any info that I haven't said. Solve the equation \[-2 \sin ^{2} \theta + \cos \theta = -2\] where \[0 \le \theta < 2 pi\] Give exact solutions.

OpenStudy (anonymous):

does that say anything else that I haven't mentioned?

OpenStudy (fibonaccichick666):

Ok, let's restart because I think you are overly stressing right now

OpenStudy (anonymous):

I am D:

OpenStudy (fibonaccichick666):

So here is your problem: \[−2sin^2θ+cosθ=−2\]

OpenStudy (fibonaccichick666):

Now, let's think about this together

OpenStudy (fibonaccichick666):

you need to find a way to pull out theta right?

OpenStudy (anonymous):

I guess so. If that's what needs to happen, I am learning and I will say yes because now I have some sort of handle.

OpenStudy (anonymous):

I guess that makes sense, because the way I've seen, I guess I just associate theta with an unknown value. Is that correct in most cases?

OpenStudy (fibonaccichick666):

yup theta is just another variable

OpenStudy (anonymous):

so this is like solving for x, only x is replaced with the symbol theta? okay got iy

OpenStudy (anonymous):

it*

OpenStudy (fibonaccichick666):

yup yup

OpenStudy (anonymous):

I'm sorry. I'm really trying ._. so yeah, we need to isolate theta

OpenStudy (fibonaccichick666):

so now, we want to only have cosines in our equation, that means we need to swap out that sin^2 with something equal to it

OpenStudy (anonymous):

(csc)?

OpenStudy (fibonaccichick666):

Any ideas on something that would be equal to \(sin^2\theta\)?

OpenStudy (anonymous):

I'm probably grasping at straws :c

OpenStudy (fibonaccichick666):

ok, not at all but tell me why you said that, I really don't knowhow you got that

OpenStudy (anonymous):

sin x = 1/csc x? ehhh

OpenStudy (fibonaccichick666):

Yes, but that's not sin^2 x

OpenStudy (fibonaccichick666):

you want specifically, sin^2 x

OpenStudy (anonymous):

ah, so it's the half angle identity

OpenStudy (fibonaccichick666):

x, theta, y, alpha whatever the variable doesn't matter

OpenStudy (fibonaccichick666):

No

OpenStudy (fibonaccichick666):

you don't have a half angle I say again

OpenStudy (fibonaccichick666):

You have to solve another for sin^2 x

OpenStudy (anonymous):

that's literally all I see that has the sin^2 (x)

OpenStudy (fibonaccichick666):

look again

OpenStudy (fibonaccichick666):

start at the beginning and just re read everything, you'll see it

OpenStudy (anonymous):

uh okay, does it have anything to do with like, sin^2(x) + cos^2(x) = 1?

OpenStudy (fibonaccichick666):

yup

OpenStudy (anonymous):

OHHHHHHHH

OpenStudy (anonymous):

is it cos^2(x) - 1 = sin^2(x) or something?

OpenStudy (fibonaccichick666):

continue

OpenStudy (anonymous):

I don't know where I'm going with it haha. It's coming slowly. It's currently stoped. Hint?

OpenStudy (fibonaccichick666):

your eq is \[−2sin^2θ+cosθ=−2 \] You want to swap something in that contains a cosine for that sin^2

OpenStudy (anonymous):

so that would be cos^2(x) + cos(x) - 1 = -2, right?

OpenStudy (fibonaccichick666):

wait, what did you forget?

OpenStudy (anonymous):

combining terms or something?

OpenStudy (fibonaccichick666):

Remember it's \[ \color{red}{(}cos^2(x) - 1 \color{red}{)}= sin^2(x)\]

OpenStudy (fibonaccichick666):

and your original is :\[\color{red}{−2}sin^2θ+cosθ=−2\]

OpenStudy (anonymous):

OH, so parenthesis, so it would be (cos^2(x) - 1) + (cos(x)) = -2, right?

OpenStudy (anonymous):

uh you highlighted the -2. What happens with the -2?

zepdrix (zepdrix):

\[\Large\rm 1-\cos^2x=\sin^2x\]Did you guys get the identity mixed up a little bit? Maybe I misread the last few comments... hmm

OpenStudy (anonymous):

OH. Is that right?

OpenStudy (fibonaccichick666):

Thank's Zep, I didn't check it

OpenStudy (anonymous):

Thank you Zep!

zepdrix (zepdrix):

Oh that wasn't the boo boo she was point out hehe XD I see

zepdrix (zepdrix):

There's another one spoon! Get it!!! :D

OpenStudy (fibonaccichick666):

And you got rid of the -2 when you put it in. It doesn't go anywhere

OpenStudy (anonymous):

oh crap! YEAH!

OpenStudy (fibonaccichick666):

So step by step for me now, start with the original, then sub in

OpenStudy (anonymous):

so it's -2(1 - cos^2(x)) + (cos(x)) = -2?

OpenStudy (anonymous):

oh haha, well I got right to the correct subbing I guess

OpenStudy (fibonaccichick666):

Now, simplify the left

OpenStudy (anonymous):

I'm sorry, time is tightening, I want to work as quick as I can. So, can I actually simplify it, because one of the terms has a power

OpenStudy (fibonaccichick666):

O.o? why is there a time limit? and simplify=get rid of some parentheses

OpenStudy (anonymous):

cos(x)+cos(2 x)-1 = -2 and there's a time limit because my family wats me off the computer soon

OpenStudy (anonymous):

wants*